# Optimus-1[Uncaptioned image]: Hybrid Multimodal Memory Empowered Agents Excel in Long-Horizon Tasks

### v1

@@ Appendix D Theory
Appendix E Benchmark Suite
E.1 Benchmark
@@
E.2 Human-level Baseline
To better demonstrate agent’s performance level in Minecraft, we hired 10 volunteers to play the game as a human-level baseline. The volunteers played the game with the same environment and settings, and every volunteer asked to perform the each task on the benchmark 10 times. Ultimately, we used the average scores of 10 volunteers as the human-level baseline. The results of the human-level baseline are shown in Table 1.
To ensure the validity of the experiment, we ensured that each volunteer had at least 20 hours of Minecraft gameplay before conducting the experiment. For each volunteer, we pay \$$\$25$ as reward.
E.3 Minecraft Agents
In this section, we summarise the differences between existing Minecraft agents. As shown in the Table 7, earlier work [1, 6, 17, 3] constructed Transformer-based policy network as agent. Recent work [34, 38, 39, 24] introduces the Multimodal Large Language Model, which empowers the agent to complete long-horizon tasks by exploiting the powerful language comprehension and planning capabilities of LLM. However, these agents lack knowledge and experience, and their performance in Minecraft is still vastly gapped from the human level. In this paper, we introduce Hybrid Multimodal Memory, which empowers Optimus-1 with hierarchical knowledge and multimodal experience. This makes Optimus-1 significantly outperform all existing agents on challenging long-horizon tasks benchmark, and exhibits near human-level performance on many tasks.
@@ E.3 Minecraft Agents
Appendix F Implementation Details
F.1 Hybrid Multimodal Memory
F.1.1 Abstracted Multimodal Experience Pool
@@
As shown in Figure 2, we first input the visual image stream captured by Optimus-1 during the execution of the task to the video buffer, which filters the image stream at a fixed frequency, which makes the length of the image stream substantially shorter. Empirically, we set the frequency of filtering to 1 second/frame, meaning that the video buffer takes one frame per second from the original image stream to compose the filtered image stream. We found that above this frequency makes the visual information redundant (too much similarity between images), and below this frequency does not preserve enough complete visual information. We then feed the filtered frames into an image buffer with a window size of 16. We dynamically compute the similarity between images in the image buffer, when a new image comes in, we compute the similarity between the new image and the most recent image, and then we remove the image with the highest similarity in order to keep the image buffer’s window size to 16. We then introduce MineCLIP [6], a pre-trained model of video-text alignment with a structure similar to CLIP [26], as our visual summariser. For a given sub-goal, it calculates the correlation between the visual content within the current memory bank and the sub-goal, and when this correlation exceeds a pre-set threshold, the frames within the memory bank are saved as the visual memories corresponding to that sub-goal. Finally, we store the visual memories with the sub goal’s textual description into the Abstracted Multimodal Experience Pool. In addition, we incorporate the environment information, agent initial state, plan from Knowledge-Guided Planner, etc. into the experience memory of the given task. In this way, we consider the history information of each sub-goal and summaries and summarise it to finally compose the multimodal experience of the given task.
@@
F.1.2 Hierarchical Directed Knowledge Graph
World knowledge is the foundation and key to an agent’s ability to perform long-horizon complex tasks. Unlike allowing agents to learn world knowledge implicitly through fine-tuning [24, 45], we propose a method for constructing and transforming hierarchical world knowledge into graph structure that allows agents to learn knowledge explicitly in a more efficient manner.
@@ F.1.2 Hierarchical Directed Knowledge Graph
Mining and crafting objects in Minecraft is a complex and large knowledge network, and has a significant impact on the ability of the agent to generate effective task planning. As shown in the Figure 2, crafting a diamond sword [Uncaptioned image] requires two diamonds [Uncaptioned image] and a wooden stick [Uncaptioned image], while mining diamonds requires an iron pickaxe [Uncaptioned image], which in turn requires additional raw materials and crafting steps. We transform this mine and craft knowledge into a graph structure, where the nodes of the graph are objects, and the nodes point to objects that can be crafted or completed by that object. With directed graph, we show that connections between objects are established, and that this knowledge can be stored and updated efficiently. For a given object, we only need to retrieve the corresponding node to extract the corresponding subgraph from the knowledge graph. Then by topological sorting, we can get the antecedents and required materials for the object, and this information is provided to the Knowledge-Guided Planner as a way to generate a more reasonable sequence of sub-goals. With Hierarchical Directed Knowledge Graph, we can significantly enhance the world knowledge of the agent in a train-free manner, as shown in the experimental results in Section 3.3.
F.2 Hybrid Multimodal Memory Driven Optimus-1
In order to implement the proposed Hybrid Multimodal Memory and to progressively increase the capacity of Optimus-1 in a self-evolution manner, we propose a non-parametric learning method named "free exploration-teacher guidance". In the free exploration phase, we randomly initialize Optimus-1’s equipment and tasks, and ask it to explore freely in a randomly set environment, and acquire world knowledge through the environmental feedback, e.g., "A stone sword [Uncaptioned image] can be crafted by using a wooden stick [Uncaptioned image] and two cobblestone [Uncaptioned image]", and then finally store it in Hierarchical Directed Knowledge Graph. At the same time, Optimus-1 stores successful and failed cases in the Abstracted Multimodal Experience Pool, which enables Optimus-1 to accumulate reference experiences in the reflection phase. In the free exploration phase, we initialize multiple Optimus-1 at the same time, and they share the same Hierarchical Directed Knowledge Graph and Abstracted Multimodal Experience Pool, thus the memory is filled up efficiently. After free exploration, Optimus-1 have basic hierarchical knowledge and multimodal experience. In the teacher guidance phase, Optimus-1 need to learn a small number of long-horizon tasks based on extra knowledge given by teacher. For example, during the free exploration phase, Optimus-1 mastered crafting stick [Uncaptioned image] and mining diamond [Uncaptioned image], but did not know that "a diamond sword [Uncaptioned image] is obtained by a stick [Uncaptioned image] and two diamonds [Uncaptioned image]". This will serve as extra knowledge to guide Optimus-1 in his attempts to complete the task of "craft diamond sword". During the teacher guidance phase, Optimus-1’s memory is further expanded and it gains the experience of executing complete long-horizon tasks.
Unlike fine-tuning, we incrementally enhance the capabilities of Optimus-1 in Minecraft without any parameters updating, in a self-evolution manner. Specifically, we initialize Optimus-1’s Hybrid Multimodal Memory to be empty, and then let Optimus-1 iterate on "free exploration-teacher guidance" learning and unseen task inference. As the number of iterations grows, Optimus-1 expands its Hybrid Multimodal Memory capacity through "free exploration-teacher guidance", allowing it to progressively master short-term to long-term tasks. The experimental results in section 3.5 demonstrate that this non-parametric learning approach empowers Optimus-1 to incrementally enhance long-horizon task performance.
F.3 Backbone of Optimus-1
@@
Optimus-1 consists of Knowledge-Guided Planner, Experience-Driven Reflector, and Action Controller. In this paper, we employ OpenAI’s GPT-4V (gpt-4-turbo) §§§https://openai.com/index/gpt-4v-system-card/ as Knowledge-Guided Planner and Experience-Driven Reflector, and STEVE-1 [17] as Action Controller. We also employ open-source models like Deepseek-VL [18] and InternLM-XComposer2-VL [5] as Knowledge-Guided Planner and Experience-Driven Reflector.
@@
All experiments were implemented on 4x NVIDIA A100 GPUs. We employ multiple Optimus-1 to perform different tasks at the same time, and this parallelized inference greatly improves our experimental efficiency. Throughout the experiment, we spent about \$$\$5,000$ to access the GPT-4V API. However, we also offer more cost-effective solutions. As shown in Figure 5, if we employ Deepseek-VL [18] or InternLM-XComposer2-VL [5] as Optimus-1’s backbone, we can get comparable performance for nearly free!
F.4 Prompt for Optimus-1
@@
Appendix G Full Experimental Results
G.1 Results on Our Benchmark
We list the results of each task on the benchmark below, with details including task name, sub-goal number of task, success rate (SR), average number of steps (AS), average time (AT), and eval times. All tasks are evaluated in Minecraft 1.16.5 Survival Mode. Note that each time Optimus-1 performs a task, we initial it with an empty initial inventory and a random start point. This makes it challenging for Optimus-1 to perform each task.
@@ G.1 Results on Our Benchmark
G.2 Results on Other Benchmark
For a more comprehensive comparison with current Minecraft Agents, we also report Optimus-1’s performances on the benchmark used by Voyager [34], MP5 [24], and DEPS [38] below. Due to the different environments and settings, agents perform tasks with varying degrees of difficulty. For example, Optimus-1 requires low-level action to perform any task in MineRL [10], and we initialize its inventory to be empty. While Voyager [34] performs tasks in Mineflayer https://github.com/PrismarineJS/mineflayer environment only through encapsulated code, MP5 [24] performs tasks in MineDOJO [6] environment only needs a specific control signal to craft tools, no low-level actions (mouse movement and click) are needed.
@@ G.2 Results on Other Benchmark
Appendix H Case Study
@@
  • Figures 10 and 11 showcase Optimus-1’s ability to make judgments based on visual signals. When Optimus-1 determines that it has completed a task (such as "kill a cow [Uncaptioned image]" in Figure 10), it will finish the current task and move on to the next one. If Optimus-1 discovers that it has not yet completed the task and the task has not failed(as shown in Figure 11), it will continue executing the task.
@@
  • Figure 12 demonstrates the importance of knowledge. For a long-horizon task such as "Mine 1 diamond [Uncaptioned image]," Optimus-1 first generates a plan based on the Hierarchical Directed Knowledge Graph. However, this plan needs to be adjusted based on the current visual signals. For example, in this figure, Optimus-1 appears in a cave, so the primary task is not to "chop a tree" but to "leave the cave" first. Only after exiting the cave can Optimus-1 proceed with the initial plan.

### v2

@@ Appendix C Theory
Appendix D Benchmark Suite
D.1 Benchmark
@@
D.2 Baselines
Existing Baseline. On the one hand, we employ GPT-3.5 and GPT-4V as baseline, which are evaluated without integrating hybrid multimodal memory modules. During the planning phase, they generate a plan for the action controller based on task prompt (and observation). During the reflection phase, they generate reflection results in a zero-shot manner. On the other hand, we compare existing SOTA Agents [50, 51] in Minecraft.
Human-level Baseline. To better demonstrate agent’s performance level in Minecraft, we hired 10 volunteers to play the game as a human-level baseline. The volunteers played the game with the same environment and settings, and every volunteer asked to perform the each task on the benchmark 10 times. Ultimately, we used the average scores of 10 volunteers as the human-level baseline. The results of the human-level baseline are shown in Table 1. To ensure the validity of the experiment, we ensured that each volunteer had at least 20 hours of Minecraft gameplay before conducting the experiment. For each volunteer, we pay \$$\$25$ as reward.
D.3 Minecraft Agents
In this section, we summarise the differences between existing Minecraft agents. As shown in the Table 7, earlier work [1, 7, 25, 3] constructed Transformer-based policy network as agent. Recent work [46, 50, 51, 32] introduces the Multimodal Large Language Model, which empowers the agent to complete long-horizon tasks by exploiting the powerful language comprehension and planning capabilities of LLM.
In the Mineflayer and Minedojo environments, agents [7, 46, 50, 32] can accomplish sub-goals by calling APIs (in the form of codes), which is a different behavioral pattern from humans. In MineRL [11], agents [1, 25, 3, 51] must generate low-level actions to perform tasks, which is more challenging to accomplish long-horizon tasks.
Moreover, existing agents lack knowledge and experience, and their performance in Minecraft is still vastly gapped from the human level. In this paper, we introduce Hybrid Multimodal Memory, which empowers Optimus-1 with hierarchical knowledge and multimodal experience. This makes Optimus-1 significantly outperform all existing agents on challenging long-horizon tasks benchmark, and exhibits near human-level performance on many tasks.
@@ D.3 Minecraft Agents
Appendix E Implementation Details
E.1 Hybrid Multimodal Memory
E.1.1 Abstracted Multimodal Experience Pool
@@
As shown in Figure 2, we first input the visual image stream to the video buffer, which filters the image stream at a fixed frequency. It makes the length of the image stream substantially shorter. Empirically, we set the frequency of filtering to 1 second/frame, meaning that the video buffer takes one frame per second from the original image stream to compose the filtered image stream. We found that above this frequency makes the visual information redundant (too much similarity between images), and below this frequency does not preserve enough complete visual information.
Then, we feed the filtered frames into an image buffer with a window size of 16. We dynamically compute the similarity between images in the image buffer, when a new image comes in, we compute the similarity between the new image and the most recent image, and then we remove the image with the highest similarity in order to keep the image buffer’s window size to 16.
Subsequently, we introduce MineCLIP [7], a pre-trained model of video-text alignment with a structure similar to CLIP [34], as our visual summariser. For a given sub-goal, it calculates the correlation between the visual content within the current memory bank and the sub-goal, and when this correlation exceeds a pre-set threshold, the frames within the memory bank are saved as the visual memories corresponding to that sub-goal. Finally, we store the visual memories with the sub goal’s textual description into the Abstracted Multimodal Experience Pool. In addition, we incorporate the environment information, agent initial state, plan from Knowledge-Guided Planner, etc. into the experience memory of the given task. In this way, we consider the history information of each sub-goal and summaries and summarise it to finally compose the multimodal experience of the given task.
@@
E.1.2 Hierarchical Directed Knowledge Graph
@@
As shown in the Figure 2, crafting a diamond sword [Uncaptioned image] requires two diamonds [Uncaptioned image] and a wooden stick [Uncaptioned image], while mining diamonds requires an iron pickaxe [Uncaptioned image], which in turn requires additional raw materials and crafting steps. We transform this mine and craft knowledge into a graph structure, where the nodes of the graph are objects, and the nodes point to objects that can be crafted or completed by that object. With directed graph, we show that connections between objects are established, and that this knowledge can be stored and updated efficiently. For a given object, we only need to retrieve the corresponding node to extract the corresponding subgraph from the knowledge graph. Then by topological sorting, we can get the antecedents and required materials for the object, and this information is provided to the Knowledge-Guided Planner as a way to generate a more reasonable sequence of sub-goals. With Hierarchical Directed Knowledge Graph, we can significantly enhance the world knowledge of the agent in a train-free manner, as shown in the experimental results in Section 3.3.
Our HDKG can be efficiently updated and expanded. When adding new nodes, the HDKG can be updated by simply merging the nodes and relationships into the graph. This method involves local linear modifications to the graph rather than altering the entire graph, making the process efficient and time-saving. For example, when M new nodes and N edges are added, the HDKG can be updated with M+N times of operations. Moreover, an HDKG containing 851 objects (nodes) requires less than 1 MB of memory. Thus, the HDKG can be efficiently updated and maintained.
E.2 Hybrid Multimodal Memory Driven Optimus-1
In order to implement the proposed Hybrid Multimodal Memory and to progressively increase the capacity of Optimus-1 in a self-evolution manner, we propose a non-parametric learning method named “free exploration-teacher guidance”.
In the free exploration phase, we randomly initialize the environment, materials, and tasks. For the task “craft a wooden pickaxe”, we provide initial materials (three planks, two sticks), and then Optimus-1 (only the action controller activated) attempts to complete the task. If the environment feedback indicates the task is successful, the knowledge {3 planks, 2 sticks → wooden pickaxe} is added to the HDKG. Note that we randomly initialize materials and their quantities, which means that the task may not always succeed. As a result, each free exploration may not acquire the corresponding knowledge, but it can record the relevant experience (whether successful or fail). In the free exploration phase, Optimus-1 learns simple atomic operations, such as crafting sticks in the Wooden Group and mining diamonds in the Diamond Group.
In the teacher guidance phase, Optimus-1 need to learn a small number of long-horizon tasks based on extra knowledge. For example, during the free exploration phase, Optimus-1 mastered crafting stick [Uncaptioned image] and mining diamond [Uncaptioned image], but did not know that “a diamond sword [Uncaptioned image] is obtained by a stick [Uncaptioned image] and two diamonds [Uncaptioned image]”. So we provide some task plans, which will serve as extra knowledge to guide Optimus-1 to complete the task of “craft diamond sword”. We built the following automated process to get the task plan needed for “free exploration”:
  • We randomly select 5 tasks for each Group (7 groups in total) that are not included in the benchmark.
  • For each selected task, we use a script to automatically obtain the crafting relationships from the Minecraft Wiki  [^4] https://minecraft.wiki/. Taking the task “craft a wooden sword” as an example, we use the script to automatically obtain the crafting relationships: 1 wooden stick, 2 planks, 1 crafting table → 1 wooden sword, 1 log → 4 planks, 2 planks → 4 sticks, 4 planks → 1 crafting table.
  • These relationships are converted into a directed acyclic graph through an automated script. By performing a topological sort, the graph can be converted into tuples of materials and their quantities: (wooden sword, 1), (crafting table, 1), (wooden stick, 1) (planks, 8), (log, 2).
  • We prompt GPT-4 to construct a plan in order from basic materials to advanced materials.
  • Finally, we get the plan: 1. Get two logs 2. Craft eight planks 3. Craft a crafting table 4. Craft a wooden stick 5. Craft a wooden sword
During the teacher guidance phase, Optimus-1’s memory is further expanded and it gains the experience of executing complete long-horizon tasks. Teacher guidance phase allows Optimus-1 to acquire advanced knowledge and learn multimodal experiences through complete long-horizon tasks.
E.3 Backbone of Optimus-1
@@
Optimus-1 consists of Knowledge-Guided Planner, Experience-Driven Reflector, and Action Controller. In this paper, we employ OpenAI’s GPT-4V (gpt-4-turbo)  [^5] https://openai.com/index/gpt-4v-system-card/ as Knowledge-Guided Planner and Experience-Driven Reflector, and STEVE-1 [25] as Action Controller. We also employ open-source models like Deepseek-VL [26] and InternLM-XComposer2-VL [6] as Knowledge-Guided Planner and Experience-Driven Reflector.
@@
All experiments were implemented on 4x NVIDIA A100 GPUs. We employ multiple Optimus-1 to perform different tasks at the same time, and this parallelized inference greatly improves our experimental efficiency. In the free exploration and teacher guidance phases, there is no need to access OpenAI’s API, and the learning process takes approximately 16 hours on 4x A100 80G GPUs. During the inference phase, it takes about 20 hours on 4x A100 80G GPUs.
Throughout the experiment, we spent about \$$\$5,000$ to access the GPT-4V API. However, we also offer more cost-effective solutions. As shown in Figure 5, if we employ Deepseek-VL [26] or InternLM-XComposer2-VL [6] as Optimus-1’s backbone, we can get comparable performance with low-cost!
E.4 Prompt for Optimus-1
@@
Appendix F Additional Experimental Results
F.1 Full Results on Our Benchmark
We list the results of each task on the benchmark below, with details including task name, sub-goal numbers, success rate (SR), average number of steps (AS), average time (AT), and eval times. All tasks are evaluated in Minecraft 1.16.5 Survival Mode. Note that each time Optimus-1 performs a task, we initial it with an empty initial inventory and a random start point. This makes it challenging for Optimus-1 to perform each task.
Moreover, in MineRL [11] environment, ’steps’ refers to the number of interactions between the agent and the environment, occurring at a frequency of 20 times per second. For example, if an agent takes 2 seconds to complete the task “chop a tree”, it interacts with the environment 40 times, resulting in a recorded steps number of 40. Experimental results show that Optimus-1’s average task completion step (AS) is significantly lower than other baselines.
@@ F.1 Full Results on Our Benchmark
F.2 Results on Other Benchmark
For a more comprehensive comparison with current Minecraft Agents, we also report Optimus-1’s performances on the benchmark used by Voyager [46], MP5 [32], and DEPS [50] below. Due to the different environments and settings, agents perform tasks with varying degrees of difficulty. For example, Optimus-1 requires low-level action to perform any task in MineRL [11], and we initialize its inventory to be empty. While Voyager [46] performs tasks in Mineflayer  [^6] https://github.com/PrismarineJS/mineflayer environment only through encapsulated code, MP5 [32] performs tasks in MineDOJO [7] environment only needs a specific control signal to craft tools, no low-level actions (mouse movement and click) are needed.
@@ F.2 Results on Other Benchmark
Appendix G Case Study
@@
  • Figures 10 and 11 showcase Optimus-1’s ability to make judgments based on visual signals. When Optimus-1 determines that it has completed a task (such as “kill a cow [Uncaptioned image]” in Figure 10), it will finish the current task and move on to the next one. If Optimus-1 discovers that it has not yet completed the task and the task has not failed(as shown in Figure 11), it will continue executing the task.
@@
  • Figure 12 demonstrates the importance of knowledge. For a long-horizon task such as “Mine 1 diamond [Uncaptioned image],” Optimus-1 first generates a plan based on the Hierarchical Directed Knowledge Graph. However, this plan needs to be adjusted based on the current visual signals. For example, in this figure, Optimus-1 appears in a cave, so the primary task is not to “chop a tree” but to “leave the cave” first. Only after exiting the cave can Optimus-1 proceed with the initial plan.